What is the minimum code distance for correcting three errors and detecting one additional error?

a) 1

b)3

c)8

d)6
Answer #1

Correct Answer: c

Explanation:

To detect 't' bit error, the minimum code distance must be 't+1'.

To correct 't' error, the minimum code distance must be '2t+1'.

So, to detect a 1-bit error, the minimum code distance must be 2.

To correct 3 error, the minimum code distance must be 2 * 3 + 1 = 6 + 1 = 7

The option a, b, and d are incorrect because minimum code distance is not enough to correct the 3-bit error.

So, option c is correct.